OpenStudy (anonymous):

Convert 1.75 g/mL to the unit g/m3. Please explain the steps :) 1.75 x 106 g/m3 1.75 x 102 g/m3 1.75 x 10-2 g/m3 1.75 x 10-6 g/m3

OpenStudy (anonymous):

ok so you have to convert mL which are cm3 to m3 and 1 cm3 is 1*10^-3 dm3 and 1*10^-6 m3 does this help?

OpenStudy (anonymous):

yes, thanks so much :)

OpenStudy (anonymous):

what did you choose for answer?

OpenStudy (anonymous):

I chose the last. Is that correct?

OpenStudy (anonymous):

nope its the first

OpenStudy (anonymous):

oh, okay, haha. this is confusing for me, sorry :P

OpenStudy (anonymous):

because the conversion is like follows 1,75 g/cm3 = 1,75 g/10^-6 m3 and when 10^-6 goes up it changes from - to + and you have 1,75 * 10^6 g/m3 its little difficult for me to explain like that cause english is not my native language but think logicaly if you have 1 g in 1 cm3 than you have to have more in 1 m3

OpenStudy (anonymous):

okay, i see, so im pretty much converting negative and positive as well which is why it wasnt -6 but +6

OpenStudy (anonymous):

when you have units you need to pay attention weather unit is up or down under /, if it is up then it stays like it is, but if it is under then you change - to + or + to -

OpenStudy (anonymous):

ok, i understand what you're saying i think

OpenStudy (anonymous):

now lets take example and it would be nice this one you already have... so lets start: 1,75 g/cm3 = 1,75 *10^3 g/ dm3 because 1 dm3 has 1*10^3 cm3 or 1 cm3 has 1*10^-3 dm3 now lets convert g to kg 1,75 * 10^3 g/dm3 = 1,75 kg/dm3 because 1 kg has 1*10^3 g or 1 g is 1*10^-3 kg do you get it now more clearly?

OpenStudy (anonymous):

yes, i do. thank you so much for your time and assistance :) i really appreciate it

OpenStudy (anonymous):

not a problem, glad to help